Grade 11Physical Chemistry

Wavelength of radiations emitted when an electron jumps from a state A to C is 2000Å and it is 6000Å when the electron jumps from state B to C. Wavelength of the radiations emitted when an electron jumps from state A to State B will beA: 2000ÅB: 3000ÅC: 4000ÅD: 6000Å

Sahil
9 Years agoGrade 11

2 Answers

Anuj
8 Years ago
From A to C, Energy=12400/3000=4.13 eV From B to C, Energy=12400/6000=2.1eV Now, if we subtract the second equation from the first, we`ll get the energy of the electron when it jumps from A to B. So, Energy=4.13-2.1 eV =2.03 ≈2eV Therefore, Wavelength= 12400/2 Angstorm =(around 6000 Angstorm)
Shounak yadav
7 Years ago
1/2000=R(1/A-1/C)
2000R/C=2000/A-1...(1) and
1/6000=R(1/B-1/C)
2000R/C=2000R/B-1/3...(2)
equating (1) and (2)
2000R/A-1=2000R/B-1/3
R(1/A-1/B)=1/200*(1-1/3)=1/2000*2/3=1/3000
Therefore, lamba=3000